# From LLMs to MLLMs to Agents: A Survey of Emerging Paradigms in Jailbreak Attacks and Defenses within LLM Ecosystem

### v2

@@ III-A1 Training Stage
Attacks during the training stage mainly refer to embedding vulnerabilities during model training, causing abnormal behavior under specific circumstances. Attackers need to modify the training data or model weights, which is usually applicable to parameter-based jailbreaks. (a) Backdoor Attack [68, 69, 70] is the intentional embedding of malicious trigger conditions during the model training phase, causing the model to produce attacker-prescribed abnormal outputs when encountering specific inputs. Typically, attackers insert a small number of samples with backdoor tags into the training data or directly modify model parameters to ensure that the model generates prohibited content or executes attack commands when inputs contain the backdoor trigger. Backdoor attacks are highly covert because the model still performs well on normal inputs and only activates the attack under specific conditions. This type of attack is especially dangerous in goal-oriented environments, particularly in automated content generation and security-sensitive tasks. (b) Distillation Attack [72, 73, 32] is a type of attack that bypasses security defenses through the knowledge distillation process. Attackers first train an unconstrained “teacher model” that is free from any security filtering or restrictions. Then, they use this teacher model to distill (train) a “student model,” transferring the teacher’s knowledge to the student. Because the teacher model lacks safety constraints, the student model may learn some behaviors that violate ethical or security requirements during distillation, thereby causing originally well-protected models to fail. (c) Tampering Attack [35, 34, 76] refers to interfering with the model’s normal behavior during training by modifying the training data or model parameters. Attackers may insert malicious samples to mislead the model into learning incorrect patterns or directly tamper with model parameters, causing abnormal behavior when facing specific inputs. The goal of tampering attacks is usually to make the model appear to perform well under normal conditions but output inappropriate results under specific inputs or environments.
@@ III-A2 Inference Stage
Attacks during the inference stage mainly refer to attackers inducing the model to output prohibited content during its usage. Such attacks generally do not require modifying model parameters and rely solely on carefully crafted inputs, making them applicable to parameter-free jailbreaks. (a) Prompt Attack [29, 27, 79] is an attack method that induces the model to bypass its built-in safety restrictions through carefully designed prompts. Attackers manipulate the words, structure, or tone in the input to force the model to generate prohibited content. For example, by using puns, metaphors, or implicit expressions, attackers can cleverly cause the model to produce answers that violate ethical or legal standards without directly touching sensitive topics. This type of attack requires no modification of the model parameters; simple input variations can alter the model’s output. (b) Adversarial Attack [81, 82, 83] refers to inducing the model to misjudge or lose safety constraints through minor input perturbations. Attackers insert subtle disturbances into the model input, which are usually imperceptible to the human eye but effectively influence the model’s inference process, causing it to output incorrect or inappropriate content. For example, adding meaningless noise to text input, changing word order, or replacing words with synonyms to bypass the model’s content filtering mechanisms. The key to this method is that the attacker precisely identifies the model’s vulnerabilities, making it unable to correctly detect and block inappropriate outputs. (c) Jailbreak Chain [87, 25, 26] is an attack method that gradually induces the model to provide prohibited information through a series of progressive prompts. Attackers typically first ask the model to explain a seemingly harmless concept and then continuously request details, gradually guiding the model into sensitive topics. Each round of prompts appears compliant until, through iterative dialogue, the attacker obtains prohibited outputs from the model. For example, an attacker may first ask, “What are chemical drugs?”, then follow with, “Which chemical drugs can be used to make powerful things?”, and further inquire, “How should these chemical drugs be used in bomb-making steps?” In this way, attackers leverage the model’s progressive reasoning to gradually break through the model’s content filtering restrictions.
• Hierarchy Based on the Impact of Attacks
@@ III-A3 Prompt Level
Prompt level jailbreaks mainly involve crafting carefully designed input prompts to bypass the model’s built-in safety constraints, thereby generating prohibited or unauthorized content. The main approaches include: (a) Prompt Disguise [89, 27, 90] aims to evade the model’s safety detection by modifying, encoding, or applying steganographic techniques to adversarial prompts. Attackers commonly use methods such as completion, replace, low-resource languages, and multi-strategy fusion to achieve this goal. Specifically, attackers may split sensitive prompts and represent key parts with whitespace characters, allowing the model to automatically complete the missing content. Alternatively, they replace some sensitive information with distracting words so that the model restores the original instruction during parsing and executes it. Additionally, attackers may translate sensitive content into low-resource languages that the model understands less well in order to avoid safety checks. More advanced attacks combine multiple strategies. For example, they first reduce prompt sensitivity by replacing parts of it, then further disguise it using low-resource language, and finally guide the model to generate the complete response via the completion mechanism. (b) Prompt Rewrite [25, 26, 41] employs indirect strategies to guide the model to first answer harmless questions and then progressively construct new prompts based on previous answers, ultimately inducing the model into sensitive domains. Moreover, some jailbreak frameworks have adaptive optimization capabilities. When an initial jailbreak attempt fails, they re-input the failed prompt for the model to rewrite and optimize or allow the prompt to be iteratively updated within the jailbreak framework, continuously improving the success rate of bypassing safety mechanisms.
@@ III-A4 Inference Level
Inference-level jailbreak primarily targets manipulating the model’s reasoning process to bypass safety mechanisms. The main methods include: (a) Scene Nesting [26, 36, 79] method constructs progressively complex contexts to subtly lead the model to reveal latent sensitive knowledge during step-by-step reasoning. This approach typically uses seemingly harmless stories, tables, or code as carriers, enabling the model to gradually touch on implicit sensitive backgrounds in the analysis process, thereby guiding the model to generate related information without explicitly requesting sensitive content. (b) Retrieval-Augmented Generation (RAG) [77, 78, 39] jailbreak method bypasses the model’s built-in knowledge barriers by integrating external knowledge bases such as Wikipedia or private data. Attackers cleverly mix real data with false information to interfere with the model’s knowledge reasoning process, making it difficult for the model to distinguish and filter out potentially harmful content during generation.
@@ III-A5 Model Level
Model-level jailbreak methods directly attack the model’s parameters, training process, or gradient information. They mainly include: (a) Gradient-based [41, 81, 32] methods manipulate the model through adversarial attacks or gradient optimization to produce unexpected outputs for specific inputs. Attackers leverage the model’s loss gradients to find the most effective input structures, thereby bypassing safety filters, or implant “trigger” patterns in inputs so that the model automatically generates jailbreak content when encountering certain characters or phrases. (b) Fine-tuning Attacks [87, 72, 83] involve additional training to make the model learn new behavioral patterns or bypass safety restrictions. Attackers implant malicious patterns in training data to induce the model to produce sensitive content when triggered by specific inputs. Furthermore, attackers may apply contrastive learning during fine-tuning to cause the model to behave inconsistently across different contexts, thereby evading safety detection mechanisms.
III-B Visibility of Attack
@@
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=10em, minimum height=3em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=rotate=0, child anchor=north, parent anchor=south, anchor=center, , where level=1 draw=child-one-line-color, text width=16em, minimum height=3em, font=, , where level=2 draw=child-two-line-color, text width=7em, minimum height=3em, font=, , where level=3 draw=child-three-line-color, minimum width=20em, font=, , where level=4 draw=child-four-line-color, minimum width=40em, font=, , [ Visibility of Attack, par, for tree=fill=root-color, [ White-box Jailbreak §$\S$ III-B1, for tree=fill=child-one-color, [ LLMs, for tree=fill=child-two-color-one, [ GCG [81]
@@
, for tree=fill=child-four-color-three ] ] ] ] [ Black-box Jailbreak §$\S$ III-B2, for tree=fill=child-one-color, [ LLMs, for tree=fill=child-two-color-one, [ PAIR [25]
@@
Figure 6: The classification of existing specific jailbreak methods can be first divided into two main categories: white-box jailbreak and black-box jailbreak. Subsequently, these methods are further categorized based on the type of system, including LLMs, MLLMs, and Agents.
As shown in Figure 6, within the LLM ecosystem, jailbreak methods are categorized into white-box and black-box attacks based on the attacker’s access to internal information. Furthermore, according to their specific targets, they are further classified into jailbreak strategies targeting LLMs, MLLMs, and Agents.
@@ III-B1 White-box Jailbreak
(a) LLMs: White-box jailbreaks targeting LLMs refer to scenarios in which the attacker has full access to the model’s internal architecture, parameters, and training details. With this level of access, researchers leverage gradient information, modify weights, or craft specific trigger samples to explore model vulnerabilities and bypass its safety mechanisms.
@@
In the early stages of research, pioneers such as Zou et al. [81] train models using multiple prompts involving different categories of sensitive content. They successfully develop a universal adversarial suffix named GCG. Experimental results show that this suffix effectively induces harmful outputs in both commercial LLMs such as ChatGPT, Bard, and Claude, and open-source models including LLaMA-2-Chat, Pythia, and Falcon. This discovery marks a significant milestone in revealing security weaknesses in content moderation for LLMs. Building on this, Liu et al. [29] introduce a hierarchical genetic algorithm known as AutoDAN, which enables the automatic generation of stealthy jailbreak prompts. This method uses manually crafted jailbreak prompts as a semantic initialization point and optimizes them through a score-based genetic mechanism. The result is a systematic identification of prompts capable of bypassing LLM safety defenses.
@@
Recently, the COLD [82] attack framework leverages an advanced constrained text generation method, COLD (Constrained Optimization with Langevin Dynamics), to conduct adversarial attacks. This framework supports automated generation of adversarial prompts under multiple constraints such as fluency, stealth, sentiment consistency, and contextual coherence. Its high level of controllability enables not only traditional suffix-style adversarial attacks but also more complex settings, such as paraphrase-constrained adversarial rewrites and stealthy insertions under positional constraints. Jones et al. [83] propose a discrete optimization algorithm called ARCA (Adversarial Rewriting with Contextual Awareness). This method jointly optimizes the input and output to achieve efficient control over adversarial text generation. ARCA demonstrates strong performance across various tasks. For example, it automatically completes derogatory statements targeting specific individuals, such as completing “Barack Obama is a legalized” to “unborn → baby murderer.” It also generates English outputs from French inputs and precisely controls the inclusion of specific names in the generated text.
@@
(b) MLLMs: For MLLMs, white-box jailbreaks involve not only manipulating textual inputs but also launching attacks through multimodal inputs such as images and audio. Researchers analyze the interaction mechanisms across modalities in fusion layers and craft special inputs to trigger security vulnerabilities during cross-modal reasoning. Bailey et al. [35] discover that VLMs are vulnerable to image hijacking during the inference phase, where adversarial images manipulate model behavior. They propose a general behavior-matching algorithm to train hijacking strategies and introduce a prompt-matching technique that allows attackers to leverage general datasets, independent of their chosen prompts, to train hijacking behaviors aligned with any user-defined textual instruction. Wang et al. [68] propose a dual-objective optimization strategy. First, adversarial image prefixes are optimized from random noise to elicit diverse harmful responses from the model even without textual input, thereby imbuing images with toxic semantics. Then, these adversarial image prefixes are combined with adversarial textual suffixes in a jointly optimized fashion to maximize the probability of the model affirmatively responding to various harmful instructions. The resulting image-text adversarial pairs are collectively referred to as the Universal Master Key (UMK). Luo et al. [69] introduce Cross-Prompt Attack (CroPA), which uses learnable prompts to update visual adversarial perturbations, weakening their misleading effects and significantly enhancing adversarial sample transferability across different prompts. Schlarmann et al. [70] present a novel framework to evaluate the vulnerability of multimodal models to adversarial visual attacks. They demonstrate that imperceptible visual perturbations ($\epsilon_{\infty}=\frac{1}{255}$) alter subtitle outputs of foundational multimodal models. This technique is exploited by malicious content providers to mislead honest users, for instance, by directing them to malicious websites or spreading disinformation. The study examines two types of attacks: targeted attacks, which manipulate the model to generate specific outputs, and untargeted attacks, which aim to degrade the quality of the model’s output.
@@
Recently, the assumption of poisoned (image, text) pairs in training data emerges as a critical concern in the study of multimodal jailbreak attacks. Based on this assumption, Tao et al. [34] replace original text captions with malicious jailbreak prompts, thereby enabling jailbreak attacks via poisoned images. Xu et al. [76] propose Shadowcast, a stealthy data poisoning attack in which poisoned samples are visually indistinguishable from benign images that match the corresponding texts, making detection extremely difficult. Experiments show that Shadowcast is highly effective under two attack settings. The first is Label Attack, which misleads the VLM to output incorrect class labels, such as misidentifying Donald Trump as Joe Biden. The second is Persuasion Attack, which exploits the VLM’s text generation ability to produce seemingly rational yet misleading narratives. For example, it describes junk food as healthy, thus misleading users. Wang et al. [72] propose a novel attack strategy called Stop Reasoning Attack, which bypasses the Chain-of-Thought (CoT) reasoning process during model inference. In Visual Question Answering (VQA) tasks, questions are input into the MLLM to obtain an answer. Normally, explicit prompt tokens are added after the question and options to guide the model into CoT reasoning. However, under the influence of the Stop Reasoning Attack, MLLMs skip the reasoning steps and directly produce an answer without any rationale, thereby undermining the intended CoT mechanism.
@@
(c) Agents: White-box jailbreak techniques targeting agents focus on dissecting their system architectures. The goal is to explore how, under the assumption that internal mechanisms are known, one can manipulate core decision-making modules such as task planning components, tool invocation interfaces, and memory retrieval systems to exert targeted influence on the LLM’s key functionalities.
@@
In early-stage studies, Chen et al. [38] propose AGENTPOISON, a novel red-teaming approach and the first backdoor attack framework for general LLM agents and RAG (Retrieval-Augmented Generation)-based agents. This method achieves covert manipulation by poisoning the agent’s long-term memory or RAG knowledge base. Specifically, AGENTPOISON adopts a constrained optimization framework to generate backdoor triggers, mapping the trigger instances into a unique embedding space to enhance backdoor effectiveness. This strategy ensures that whenever user inputs contain the optimized backdoor triggers, malicious demonstrations are likely to be retrieved from the poisoned memory or knowledge base, influencing the model’s output.
@@
Subsequently, Yu et al. [87] investigate the security of multi-agent networks from a topological perspective, aiming to identify which topological properties help build safer networks. To this end, they propose a universal framework named NetSafe, which integrates various LLM-based agent frameworks through iterative RelCom interactions, laying a foundation for generalized research on topological safety. In the process, the authors identify critical phenomena triggered when multi-agent networks are subjected to misinformation, bias, or harmful content. These phenomena are termed Agent Hallucination and Aggregation Safety, describing the adverse impacts such attacks have on network stability. Tan et al. [73] explore a novel vulnerability in MLLM-based societies: indirect propagation of malicious content. Unlike direct generation of harmful outputs by MLLMs, their study demonstrates how a single MLLM agent is subtly manipulated to craft specific prompts, thereby inducing other MLLM agents in the network to generate harmful content.

### v3

@@ III-A1 Training Stage
Attacks during the training stage typically involve the injection of vulnerabilities or backdoors into the model, which can lead to abnormal or malicious behavior when triggered by specific inputs. Such attacks require tampering with the training data or model parameters and are generally associated with parameter-based jailbreaks. Specifically, these training-stage attacks can be broadly categorized into the following three types: (a) Backdoor Attack [67, 68, 69] is the intentional embedding of malicious trigger conditions during the model training phase, causing the model to produce attacker-prescribed abnormal outputs when encountering specific inputs. Typically, attackers either inject a small number of samples containing backdoor triggers into the training data or directly manipulate the model parameters. The goal is to ensure that, once deployed, the model produces prohibited content or executes malicious commands whenever the input contains the predefined trigger. (b) Distillation Attack [71, 72, 32] is a type of attack that bypasses security defenses through the knowledge distillation. Attackers first train an unconstrained “teacher model” that is free from any security filtering or restrictions. Then, they use this teacher model to distill (train) a “student model,” transferring the teacher’s knowledge to the student. Since the teacher model lacks safety constraints, the student model may learn some behaviors that violate ethical or security requirements during distillation, thereby rendering the student model’s original defense mechanisms ineffective. (c) Tampering Attack [35, 34, 75] refers to interfering with the model’s normal behavior during training by modifying the training data or model parameters. The goal of tampering attacks is usually to make the model appear to perform well under normal conditions but output inappropriate results under specific inputs or environments.
@@ III-A2 Inference Stage
Attacks during the inference stage mainly refer to attackers inducing the model to output prohibited content during its usage. Such attacks generally do not require modifying model parameters and rely solely on carefully crafted inputs, making them applicable to parameter-free jailbreaks. In particular, these inference-stage attacks can be broadly classified into the following three types: (a) Prompt Attack [29, 27, 78] is an attack method that induces the model to bypass its built-in safety restrictions through carefully designed prompts. Attackers manipulate the words, structure, or tone in the input to force the model to generate prohibited content. For example, by using puns, metaphors, or implicit expressions, attackers can cleverly cause the model to produce answers that violate ethical or legal standards without directly touching sensitive topics. (b) Adversarial Attack [80, 81, 82] refers to inducing the model to misjudge or lose safety constraints through minor input perturbations. Attackers insert subtle disturbances into the model input, which are usually imperceptible to the human eye but effectively influence the model’s inference process, causing it to output incorrect or inappropriate content. For example, adding meaningless noise to text input, changing word order, or replacing words with synonyms to bypass the model’s content filtering mechanisms. (c) Jailbreak Chain [86, 25, 26] is an attack method that gradually induces the model to provide prohibited information through a series of progressive prompts. Attackers typically initiate the interaction by requesting an explanation of a seemingly benign concept. They then issue a sequence of prompts that appear legitimate, but incrementally shift the discussion toward sensitive content. Through this iterative dialogue, the attacker is able to bypass safeguards and ultimately get restricted responses from the model. For example, an attacker may first ask, “What are chemical drugs?”, then follow with, “Which chemical drugs can be used to make powerful things?”, and further inquire, “How should these chemical drugs be used in bomb-making steps?” In this way, attackers leverage the model’s progressive reasoning to gradually break through the model’s content filtering restrictions.
• Intervention Level of Attack
@@ III-A3 Prompt Level
Prompt level jailbreaks mainly involve crafting carefully designed input prompts to bypass the model’s built-in safety constraints, thereby generating prohibited or unauthorized content. The main methods include: (a) Prompt Disguise [88, 27, 89] aims to evade the model’s safety detection by modifying, encoding, or applying steganographic techniques to adversarial prompts. Attackers commonly use methods such as completion, replace, low-resource languages, and multi-strategy fusion to achieve this goal. Specifically, attackers may split sensitive prompts and represent key parts with whitespace characters, allowing the model to automatically complete the missing content. Alternatively, they replace some sensitive information with distracting words so that the model restores the original instruction during parsing and executes it. Additionally, attackers may translate sensitive content into low-resource languages that the model understands less well in order to avoid safety checks. In general, advanced attacks combine multiple strategies. For example, Mao et al. [89] first reduce prompt sensitivity by replacing parts of it, then further disguise it using low-resource language, and finally guide the model to generate the complete response via the completion mechanism. (b) Prompt Rewrite [25, 26, 41] employs indirect strategies to guide the model to first answer harmless questions and then progressively construct new prompts based on previous answers, ultimately inducing the model into sensitive domains. Moreover, some jailbreak frameworks have adaptive optimization capabilities. When an initial jailbreak attempt fails, they re-input the failed prompt for the model to rewrite and optimize or allow the prompt to be iteratively updated within the jailbreak framework, continuously improving the success rate of bypassing safety mechanisms.
@@ III-A4 Inference Level
Inference-level jailbreak primarily targets manipulating the model’s reasoning process to bypass safety mechanisms. The main methods include: (a) Scene Nesting [26, 36, 78] method constructs progressively complex contexts to subtly lead the model to reveal latent sensitive knowledge during step-by-step reasoning. This approach typically uses seemingly harmless stories, tables, or code as carriers, enabling the model to gradually touch on implicit sensitive backgrounds in the analysis process, thereby guiding the model to generate harmful response without explicitly requesting sensitive content. (b) Retrieval-Augmented Generation (RAG) [76, 77, 39] jailbreak method bypasses the model’s built-in knowledge barriers by integrating external knowledge bases such as Wikipedia or private data. Attackers cleverly mix real data with false information to interfere with the model’s knowledge reasoning process, making it difficult for the model to distinguish and filter out potentially harmful content during generation.
@@ III-A5 Model Level
Model-level jailbreak methods directly attack the model’s parameters, training process, or gradient information. They mainly include: (a) Gradient-based [41, 80, 32] methods manipulate the model through adversarial attacks or gradient optimization to produce unexpected outputs for specific inputs. Attackers leverage the model’s loss gradients to find the most effective input structures, thereby bypassing safety filters, or implant “trigger” patterns in inputs so that the model automatically generates jailbreak content when encountering certain characters or phrases. (b) Fine-tuning Attacks [86, 71, 82] involve additional training to make the model learn new behavioral patterns. Attackers implant malicious patterns in training data to induce the model to produce sensitive content when triggered by specific inputs. Furthermore, attackers may apply contrastive learning during fine-tuning to cause the model to behave inconsistently across different contexts, thereby evading safety detection mechanisms.
III-B Permissions of Attacker
@@
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=5em, minimum height=3em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=rotate=0, child anchor=north, parent anchor=south, anchor=center, , where level=1 draw=child-one-line-color, text width=16em, minimum height=3em, font=, , where level=2 draw=child-two-line-color, text width=7em, minimum height=3em, font=, , where level=3 draw=child-three-line-color, minimum width=30em, font=, , where level=4 draw=child-four-line-color, minimum width=50em, font=, , [ Permissions of Attacker, par, for tree=fill=root-color, [ White-box Jailbreak $\S$ III-B1, for tree=fill=child-one-color, [ LLMs, for tree=fill=child-two-color-one, [ GCG [80]
@@
, for tree=fill=child-four-color-three ] ] ] ] [ Black-box Jailbreak $\S$ III-B2, for tree=fill=child-one-color, [ LLMs, for tree=fill=child-two-color-one, [ PAIR [25]
@@
Figure 6: The existing jailbreak methods can be divided into two main categories from the perspective of attacker permissions: white-box jailbreak and black-box attack. Subsequently, these methods are further categorized based on the type of system, including LLMs, MLLMs, and Agents.
As shown in Figure 6, jailbreak methods within the LLM ecosystem are first categorized into white-box and black-box attacks, depending on the attacker’s access to internal model information. Based on their specific targets, these methods are further classified into jailbreak strategies aimed at LLMs, MLLMs, and Agents.
@@ III-B1 White-box Jailbreak
(a) LLMs: White-box jailbreaks targeting LLMs refer to scenarios in which the attacker has full access to the model’s internal architecture, parameters and training details. With this level of access, researchers leverage gradient information, modify weights, or craft specific trigger samples to explore model vulnerabilities and bypass its safety mechanisms.
@@
In the early stages of research, pioneers like Zou et al. [80] train models using multiple prompts across different sensitive content categories. They develop a universal adversarial suffix, GCG, which successfully induces harmful outputs in both commercial LLMs and open-source LLMs. This highlights security weaknesses in LLM content moderation. Building on this, Liu et al. [29] introduce AutoDAN, a hierarchical genetic algorithm that generates stealthy jailbreak prompts by optimizing manually crafted prompts through a score-based mechanism. This approach systematically identifies prompts that can bypass LLM safety defenses.
@@
Recently, the COLD attack framework [81] uses the Constrained Optimization with Langevin Dynamics (COLD) method to automate adversarial prompt generation under various constraints, including fluency, stealth, sentiment consistency, and contextual coherence. This enables both traditional suffix-style attacks and more complex ones, such as paraphrase-constrained rewrites and stealthy insertions. Jones et al. [82] introduce Adversarial Rewriting with Contextual Awareness (ARCA), a discrete optimization algorithm that optimizes both input and output for precise control over adversarial text generation. ARCA excels in tasks like completing derogatory statements (e.g., “Barack Obama is a legalized” to “unborn → baby murderer”), generating English from French, and controlling name inclusion in text.
@@
(b) MLLMs: For MLLMs, white-box jailbreaks involve not only manipulating textual inputs but also launching attacks through multimodal inputs such as images and audio. Researchers analyze the interaction mechanisms across modalities in fusion layers and craft special inputs to trigger security vulnerabilities during cross-modal reasoning. Bailey et al. [35] find that VLMs are vulnerable to image hijacking during inference, where adversarial images manipulate model behavior. They propose a behavior-matching algorithm and a prompt-matching technique, allowing attackers to train hijacking behaviors with general datasets, independent of the prompt. Wang et al. [67] introduce a dual-objective optimization strategy: optimizing adversarial image prefixes from random noise to induce harmful responses, then combining them with adversarial textual suffixes to maximize harmful model responses. These image-text pairs are called the Universal Master Key (UMK). Luo et al. [68] propose Cross-Prompt Attack (CroPA), which updates visual adversarial perturbations using learnable prompts, improving adversarial sample transferability. Schlarmann et al. [69] present a framework to evaluate the vulnerability of multimodal models to adversarial visual attacks, showing that imperceptible visual perturbations can alter subtitle outputs. These attacks, including targeted and untargeted types, can mislead users, directing them to malicious sites or spreading disinformation.
@@
Recently, the assumption of poisoned (image, text) pairs in training data has become a critical concern in multimodal jailbreak attacks. Tao et al. [34] exploit this by replacing original text captions with malicious jailbreak prompts, enabling attacks via poisoned images. Xu et al. [75] introduce Shadowcast, a stealthy data poisoning attack where poisoned samples are visually indistinguishable from benign images, making detection difficult. Shadowcast works in two settings: the Label Attack, which misleads VLMs into outputting incorrect labels (e.g., misidentifying Donald Trump as Joe Biden), and the Persuasion Attack, which manipulates text generation to produce misleading narratives (e.g., presenting junk food as healthy). Wang et al. [71] propose the Stop Reasoning Attack, which bypasses the Chain-of-Thought (CoT) reasoning process in Visual Question Answering (VQA) tasks. This attack causes the model to skip reasoning steps, producing answers without rationale and undermining the CoT mechanism.
@@
(c) Agents: White-box jailbreak techniques targeting agents focus on analyzing their internal system architecture. Assuming knowledge of the underlying mechanisms, these approaches aim to strategically manipulate core decision-making modules—such as task planning components, tool invocation interfaces, and memory retrieval systems—to precisely influence and control the key functionalities of agents.
@@
In early studies, Chen et al. [38] introduce AGENTPOISON, a novel red-teaming method and the first backdoor attack framework for general LLMs and RAG-based agents. The approach manipulates the agent’s long-term memory or knowledge base by poisoning it. AGENTPOISON uses a constrained optimization framework to generate backdoor triggers, mapping them into a unique embedding space to improve their effectiveness. This ensures that when the user input contains the optimized trigger, the malicious demonstration can be retrieved from the poisoned memory, thereby influencing the model’s output.
@@
Subsequently, Yu et al. [86] investigate the security of multi-agent networks from a topological perspective, aiming to identify properties that contribute to safer networks. They propose NetSafe, a universal framework that integrates various LLM-based agent systems through iterative RelCom interactions, establishing a foundation for studying topological safety. They identify critical phenomena, such as Agent Hallucination and Aggregation Safety, which describe the negative impact of misinformation, bias, or harmful content on network stability. Tan et al. [72] explore a novel vulnerability in MLLM-based societies: indirect propagation of malicious content. Unlike direct generation of harmful outputs by MLLMs, their study demonstrates how a single MLLM agent is subtly manipulated to craft specific prompts, thereby inducing other MLLM agents in the network to generate harmful content.